A 104 m3 thoroughly mixed pond has a water inflow and outflow of 5 m3/h. The inflow water contains 0.01 mol/m3 of chemical. Chemical is also discharged directly into the pond at a rate of 0.1 mol/h. There is no reaction, volatilization, or other losses of the chemical; it all leaves in the outflow water. (i) What is the concentration (C) in the outflow water?
Answer #1

Since all chemical leaves the water, the inflow rate will be equal to outflowt rate.

Inflow rate = 5(m3/h)*0.01(mol/m3) + 0.1(mol/h) = 0.15 mol/h

let outflow water conc. be C

Outflow rate = 5(m3/h)*C(mol/m3) = 0.15 mol/h

C = 0.03 mol/m3

Please upvote if you like the answer. Thanks